two camels pass each other in the desert, going in oppisite directions. the rate of one camel is 3 km/h faster than the rate of the other. four hours later, the camels are 68 km apart. find the rate of the faster camel.
Answer by stanbon(75887)

two camels pass each other in the desert, going in opposite directions. the rate of one camel is 3 km/h faster than the rate of the other. four hours later, the camels are 68 km apart. find the rate of the faster camel.
---------------------------------
slower camel rate = x
faster camel rate = x+3
-------------------------------
Their separation rate is 2x+3.
Equation:
dist = rate *time
68 = (2x+3)4
17 = 2x+3
x = 7
----
rate of faster camel = x+3 = 10 km/h
==========================================
Cheers,
Stan H.